OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. can some one happy

OpenStudy (january123):

Did you have this question before??

OpenStudy (january123):

P = the profit they make n = the number of flowers they sell m = the mark-up of the flowers c = the cost of a flower Then P = mcn - cn Rearrange to find n P = n(mc - c) n = P / (mc - c) Substituting values n = 500 / (1.4*0.5 - 0.5) n = 2500 They must sell 2500 flowers to make a profit of $500.

OpenStudy (anonymous):

no so that it?

OpenStudy (january123):

so they sell a flower at 140% from the original price that would be 0.50$......100% x$.............140% x= 0.50*140/100=0.70$ a flower to make 500$ you shoul sell 500/0.70= 714 flowers because they sell a flower at its original 0.50$ price plus 40% from that, meaning 100%+40%

OpenStudy (anonymous):

so that the answer?

OpenStudy (january123):

I guess so! idk

OpenStudy (anonymous):

ok thxs